# Compositional Generalization Requires More Than Disentangled Representations

### v1

@@ Abstract
Composition—the ability to generate myriad variations from finite means—is believed to underlie powerful generalization. However, compositional generalization remains a key challenge for deep learning. A widely held assumption is that learning disentangled (factorized) representations naturally supports this kind of extrapolation. Yet, empirical results are mixed, with many generative models failing to recognize and compose factors to generate out-of-distribution (OOD) samples. In this work, we investigate a controlled 2D Gaussian “bump” generation task, demonstrating that standard generative architectures fail in OOD regions when training with partial data, even when supplied with fully disentangled $(x,y)$ coordinates, re-entangling them through subsequent layers. By examining the model’s learned kernels and manifold geometry, we show that this failure reflects a “memorization” strategy for generation through the superposition of training data rather than by combining the true factorized features. We show that models forced—through architectural modifications with regularization or curated training data—to create disentangled representations in the full-dimensional representational (pixel) space can be highly data-efficient and effective at learning to compose in OOD regions. These findings underscore that bottlenecks with factorized/disentangled representations in an abstract representation are insufficient: the model must actively maintain or induce factorization directly in the representational space in order to achieve robust compositional generalization.
@@ 1 Introduction
In this paper, we develop a mechanistic perspective on why simply factorizing an intermediate (or input) representation often falls short of enabling robust out-of-distribution (OOD) generalization. Concretely, we focus on a synthetic 2D Gaussian “bump” generation task, where a network learns to decode $(x,y)$ into a spatial image. We provide a detailed empirical investigation into why disentangled representations often fail to achieve robust compositional generalization. Specifically:
  • From a manifold-based viewpoint, we show that while the input or bottleneck layer may be factorized, subsequent layers can “warp” and remix factors, leading to poor out-of-distribution (OOD) extrapolation.
  • Using a kernel-based perspective, we show that networks failing to compose have overwritten their factorized inputs by “memorizing” training data, rather than combining independent factors, and simply superpose memorized states when attempting OOD generalization.
  • We use a Jacobian-based metric tensor to study the representational manifold, demonstrating how factorization erodes layer-by-layer in standard CNN decoders.
  • Finally, we show that enforcing factorization directly in the output (pixel) space—through architectural constraints or curated datasets—enables the model to overcome memorization strategies and learn genuinely compositional rules. This approach leads to data‐efficient, out‐of‐distribution generalization in cases where factorized latents alone would otherwise fail.
In addition, we provide in the Appendix A a general framework for interpreting compositionality and factorization that further clarifies these observations. Overall, these results caution against the assumption that a factorized bottleneck automatically confers compositional extrapolation, and they point toward more comprehensive design strategies for achieving genuinely compositional neural models.
@@ 2 Related Work
Factorization and compositional generalization have been widely investigated in various architectures (Zhao et al., 2018; Higgins et al., 2017b; Burgess et al., 2018; Montero et al., 2021; Xu et al., 2022; Okawa et al., 2023; Wiedemer et al., 2023; Lippl & Stachenfeld, 2024; Liang et al., 2024b). For instance, Xu et al. (2022) introduced an evaluation protocol to assess compositional generalization in unsupervised representation learning, discovering that even well-disentangled representations do not guarantee improved out-of-distribution (OOD) behavior. Similarly, Montero et al. (2021) concluded that a model’s capacity to compose novel factor combinations can be largely decoupled from its degree of disentanglement. These observations temper the once-common assumption that factorized latents automatically yield systematic compositional generalization. While many works study how $\beta$-VAEs and related methods learn disentangled representations (Burgess et al., 2018; Higgins et al., 2017a), they typically do not further explore the nuanced roles these representations play in compositional extrapolation—particularly in the presence of mixed discrete and continuous features. Consequently, a gap remains in understanding how—and to what extent—models that appear to disentangle latent factors can also robustly compose them under distribution shifts. Our work aims to illuminate this gap by investigating the detailed mechanics of factorization and compositionality within neural decoders, providing a complementary perspective to prior disentanglement benchmarks and broader generative modeling approaches.
3 Why Disentangled Bottlenecks Often Fail Compositional Generalization
We present a toy 2D Gaussian “bump” generation task that highlights a core phenomenon: even with explicitly disentangled inputs or bottleneck, standard feedforward networks can fail to generalize compositionally.
@@ 3.1 Toy Example: 2D Gaussian Bump Generation
Figure 1: Various disentangled input encodings fail to support compositional generalization. (a) shows the experimental setup for the 2D Gaussian bump toy experiment, where the scalar $(x,y)$ coordinate pair is encoded into disentangled representation (population-based vs. rate-based encodings), which is then fed to a decoder-only architecture to generate a $N\times N$ grayscale image of a single 2D Gaussian bump centered at that corresponding $(x,y)$ location. The training dataset excludes all images that contain 2D Gaussian bumps centered within the red-shaded OOD region in the center of the image field. (b-c) shows the MSE error contour plots and sample generated OOD images of bump-based and rate-based encoding of the $x$ and $y$ input, the compositional OOD region is marked with by the red-dashed bounding box and the ground truth bump location is marked by a red cross. For a non-compositional network trained with bump-encoded inputs, (d) demonstrates that it learns to “superpose” seen ID training data when asked to compositionally generalize, and (e) shows the agreement between a theoretical binary factorized kernel with the similarity matrix (computed based on the pixel overlap) between the model’s ID and OOD generated samples.
@@
We consider a generative model that receives a 2D coordinate $(x,y)$ within some bounding box (e.g. $[0,N]^{2}$) and outputs a $N\times N$ grayscale image containing a Gaussian “bump” at location $(x,y)$. The in-distribution training set might cover an “O”-shaped region of $(x,y)$ with a held-out OOD region in the center of the training distribution (Fig. 1(a)).
@@
In this work, we consider a scenario where an autoencoder’s latent bottleneck is explicitly regularized to learn a disentangled representation, and ask whether a decoder can truly leverage that factorization for compositional generalization. Concretely, we focus on a CNN-based “decoder‐only” architecture that maps disentangled latent inputs to image outputs. We also experiment with MLP decoders and provide those results in Appendix E.3, observing similar qualitative findings.
@@
Previous work has shown evidence indicating that vector-based representations are key to compositional generalization (Yang et al., 2023). Inspired by different forms of stimuli coding mechanisms within the biological brains (Georgopoulos et al., 1986; O’Keefe & Nadel, 1978; Shadlen & Newsome, 1998), we test various input encodings:
@@
    Rate-based coding, a scalar $(x,y)$ analog representation.
In each case, the input representation is trivially factorized w.r.t. $x$ and $y$ (meaning $x$ and $y$ are independently encoded by separate neurons). We then feed these inputs into a CNN that upsamples or decodes into an $N\times N$ image. For the remainder of the manuscript, we mostly compare and contrast models trained with the local-bump input encoding ($x$ or $y$ inputs embedded as the center of a 1D Gaussian bump, a vector of length $N$) and scalar input encoding ($x$ or $y$ inputs as scalar values).
3.2 Model fails to compositionally generalize for all disentangled encoding types.
Despite the factorized or “disentangled” input, the model does not automatically learn to generalize beyond the ID region. In many runs, the CNN effectively memorizes the training distribution and fails to correctly place the bump in the OOD region. Fig. 1(b-c) show the MSE error landscapes of generated images (from ground truth images) as well as their corresponding generated OOD images for models trained with the two different input encodings. While networks trained with different input encodings demonstrate different degrees of “local” generalization, all networks seem to consistently generate bizarre artifacts such as multiple bumps, which lend insights into the learned mode of “compositional” generalization of the network in relation to memorization. Hence, simply providing factorized inputs, regardless of the form of encodings, does not suffice. The deeper layers can re-entangle $x$ and $y$ or rely on memorized local features, leading to poor compositional extrapolation. While not shown, we have also experimentally verified that models trained with 1-hot and positional input encodings also fail to compositionally generalize, despite their factorized structure. These findings confirm that factorization alone, regardless of the input/bottleneck encoding formats, is not sufficient for compositional generalization.
@@
From Fig. 1(b-c), we observe consistent OOD generation pattern across the bump-based and rate-based input encodings. We note that similar generating patterns have been reported for CNN-based Diffusion model (Liang et al., 2024a). The consistency of these patterns across input encoding formats and architectures indicate a shared underlying mechanism of model learning and generalization that we further characterize here.
Figure 2: Failure of compositional generalization despite factorized input: input memorization undoes factorization. (a-b) show the linear probe metrics, and (c-d) show the factorization score (Eq. 7), and (e-f) show the volume metric $\log_{10}(dv)$, all as a function of layer depth, for models trained with bump-based vs. rate-based input encodings. The linear probe metric is defined as the $R^{2}$ scores of fitting two linear classifiers w.r.t. $x$ and $y$.
@@
where $\kappa_{x}\left(x_{i},x_{j}\right)=\kappa_{y}\left(y_{i},y_{j}\right)=\delta_{i,j}$ is given by the Kronecker delta function. With normalization, the Gram matrix element is given by $\tilde{K}_{i,j}=\frac{K_{\text{binary}}(i,j)}{\sum_{j^{\prime}}K_{\text{binary}}\left(i,j^{\prime}\right)}$. In Fig. 1(e), we show the agreement between the binary factorized kernel (top left) and the similarity matrix between the OOD and ID generated samples. This agreement shows that the factorized input structure is allowing the model to map $x$ and $y$ coordinates independently to ID samples. Nonetheless, the model fails to generalize OOD because it has not learned disentangled factors associated with either. A full kernel mapping between ID and OOD generated samples can be found in Appendix 8.
In a set of follow-up experiments on MNIST image rotation, we observe that this same “superposition” of memorized activation patterns—drawn from related in-distribution data—emerges as a general OOD generalization strategy in neural networks that have learned to memorize ID data. Crucially, this phenomenon applies not only to compositional tasks but also to single‐dimension interpolation and extrapolation. We detail these results in Appendix E.2.
@@
Why is disentangled input insufficient for learning disentangled factors of variation? Here we seek deeper mechanistic insights into model’s learned representations along the network layers. Specifically, we track how strongly $x$ and $y$ remain disentangled across layers. We find that while the input layer is trivially factorized, subsequent convolutional layers produce mixed features with large cross-terms, particularly near the final decoding. This “manifold warping” effectively destroys compositional structure, leaving the network reliant on patterns that do not extend to unseen $(x,y)$ combinations.
@@
$\mathrm{d}v(x,y)=\sqrt{\det\bigl{(}g(x,y)\bigr{)}}\mathrm{d}x\,\mathrm{d}y$ (6)
for representation output at each of the network layer. The metric visualizations $\log_{10}(dv)$ are compared between networks with bump-based vs. rate-based encodings in Fig. 2(e) and (f). Visually, we observe significant distortion as a function of network depth, especially in the OOD region. To accompany the distortion analysis, we plot the linear probe metric (linear classifier $R^{2}$ w.r.t $x$ and $y$) as well as the factorization metric proposed based $g$ in Eq. (7) as a function of network depth in Fig. 2(a-d). Here the factorization metric is defined as
@@
where $g_{ij}$’s are the matrix elements of the $2\times 2$ matrix $g$. We note that both metrics decay relatively gracefully over the layer depth for both networks with bump-based and rate-based input encodings. Intriguingly, shallow networks with rate-based encoding experiences a sharper drop in factorization. This is likely due to the higher demand for dimension expansion with the rate-based 2-neuron input, which leads to higher distortion when allowed fewer network layers to accommodate the expansion. The volume metric visualization as well as the linear probe metric as a function of layer depth for different network depths are shown in Fig. 11 and Fig. 12 (Fig. 13 and Fig. 14 for MLPs) as well as generated OOD images in Fig. 10 in Appendix E.3.

### v2

# Compositional Generalization via Forced Rendering of Disentangled Latents

@@ Abstract
Composition—the ability to generate myriad variations from finite means—is believed to underlie powerful generalization. However, compositional generalization remains a key challenge for deep learning. A widely held assumption is that learning disentangled (factorized) representations naturally supports this kind of extrapolation. Yet, empirical results are mixed, with many generative models failing to recognize and compose factors to generate out-of-distribution (OOD) samples. In this work, we investigate a controlled 2D Gaussian “bump” generation task with fully disentangled $(x,y)$ inputs, demonstrating that standard generative architectures still fail in OOD regions when training with partial data, by re-entangling latent representations in subsequent layers. By examining the model’s learned kernels and manifold geometry, we show that this failure reflects a “memorization” strategy for generation via data superposition rather than via composition of the true factorized features. We show that when models are forced—through architectural modifications with regularization or curated training data—to render the disentangled latents into the full-dimensional representational (pixel) space, they can be highly data-efficient and effective at composing in OOD regions. These findings underscore that disentangled latents in an abstract representation are insufficient and show that if models can represent disentangled factors directly in the output representational space, it can achieve robust compositional generalization.
@@ 1 Introduction
In this paper, we develop a mechanistic perspective on why even fully disentangling the intermediate (or input) representations is often insufficient to enable robust out-of-distribution (OOD) generalization. Concretely, we focus on a synthetic 2D Gaussian “bump” generation task, where a network learns to decode given $(x,y)$ coordinates into a spatial image. We provide a detailed empirical investigation into why disentangled representations often fail to achieve robust compositional generalization. Specifically, we show that:
  • From a representational manifold viewpoint, even if the input or bottleneck (latent) layer are fully disentangled, subsequent layers usually “warp” and remix factors, leading to poor out-of-distribution (OOD) extrapolation.
  • From a kernel-based perspective, networks failing to compose have overwritten their disentangled inputs by “memorizing” training data, rather than combining independent factors; they simply superpose memorized states when attempting OOD generalization.
  • Using a Jacobian-based metric tensor to study the representational manifold, there is a layer-by-layer erosion of network disentanglement in standard CNN decoders.
  • Finally, forcing disentangled inputs to render into a representation that matches the output (pixel) space — via architectural constraints or curated datasets—enables the model to overcome memorization strategies and learn genuinely compositional rules. This approach leads to data‐efficient, out‐of‐distribution generalization in cases where factorized latents alone fail.
In addition, we provide in Appendix A a general framework for interpreting compositionality and factorization that further clarifies these observations. Overall, these results caution against the assumption that a factorized bottleneck automatically confers compositional extrapolation, and they point toward more comprehensive design strategies, such as disentangled processing and data curriculum that enforce the rendering of disentangled latents, to achieve genuinely compositional neural models.  [^2] Code available at github.com/qiyaoliang/DisentangledCompGen
@@ 2 Related Work
Factorization and compositional generalization have been widely investigated in various architectures (Zhao et al., 2018; Higgins et al., 2017b; Burgess et al., 2018; Montero et al., 2021; Xu et al., 2022; Okawa et al., 2023; Wiedemer et al., 2023; Lippl & Stachenfeld, 2024; Liang et al., 2024b). While many works study how $\beta$-VAEs and related methods can learn disentangled representations (Burgess et al., 2018; Higgins et al., 2017a), they typically do not explore whether these representations solve the problem of compositional extrapolation—particularly in the presence of mixed discrete and continuous features. However, Xu et al. (2022) introduced an evaluation protocol to assess compositional generalization in unsupervised representation learning, discovering that even well-disentangled representations do not guarantee improved out-of-distribution (OOD) behavior. Similarly, Montero et al. (2021, 2022) concluded that a model’s capacity to compose novel factor combinations can be largely decoupled from its degree of disentanglement; and (Lippl & Stachenfeld, 2024) derived and demonstrated numerous failure modes in compositional generalization with disentangled latents. These observations temper the once-common assumption that factorized latents automatically yield systematic compositional generalization. Consequently, a gap remains in understanding how—and to what extent—models that appear to disentangle latent factors can also robustly compose them under large distribution shifts, and how to encourage such generalization. Our work aims to rigorously illuminate this gap in the context of a controlled task that enables characterization of computational generalization performance starting from exactly disentangled latents. Generalization is tested under large distributional shifts, and we investigate the detailed mechanisms of compositionality by the deep neural network decoders, providing a complementary perspective to prior disentanglement benchmarks and broader generative modeling approaches.
3 Why Disentangled Latents Often Fail Compositional Generalization
We present a toy 2D Gaussian “bump” generation task to highlight and analyze a core phenomenon: even with explicitly disentangled inputs or bottleneck, standard feedforward networks can fail to generalize compositionally.
@@ 3.1 Toy example: 2D Gaussian bump generation
Figure 1: Various disentangled input/latent encodings fail to support compositional generalization. (a) shows the experimental setup for the 2D Gaussian bump toy experiment, where the scalar $(x,y)$ coordinate pair is encoded into disentangled representation (population-based vs. ramp-based encodings), which is then fed to a decoder-only architecture to generate a $N\times N$ grayscale image of a single 2D Gaussian bump centered at that corresponding $(x,y)$ location. The training dataset excludes all images that contain 2D Gaussian bumps centered within the red-shaded OOD region in the center of the image field. (b-c) shows the MSE error contour plots and sample generated ID/OOD images of bump-based and ramp-based encoding of the $x$ and $y$ input, the compositional OOD region is marked with by the red-dashed bounding box and the ground truth bump location is marked by a red cross. For a non-compositional network trained with bump-encoded inputs, (d) demonstrates that it learns to “superpose” seen ID training data when asked to compositionally generalize, and (e) shows the agreement between a theoretical binary factorized kernel with the similarity matrix (computed based on the pixel overlap) between the model’s ID and OOD generated samples.
@@
We consider a generative model that must output an $N\times N$ grayscale image containing a Gaussian “bump” at a specified 2D $(x,y)$ location within some bounding box (e.g. $[0,N]^{2}$). The training set covers a square-donut shaped region of $(x,y)$ with a large OOD region in the center of the training distribution (Fig. 1(a)); alternatively, the OOD region is an equivalent area in the bottom left corner of image. In both cases, the model sees all $x$ and $y$ values in training but many combinations are held out.
@@
To distinguish the behavior of the decoder from the encoding process that must learn (and may fail to fully disentangle) its latent representation, we explicitly construct a fully disentangled latent representation and ask whether the downstream decoder can leverage that factorization for compositional generalization. Concretely, we focus on a CNN-based “decoder‐only” architecture that maps disentangled latent inputs to image outputs. We also experiment with MLP decoders and provide those results in Appendix F.3, observing qualitatively similar findings.
@@
Previous work suggests that vector-based representations may be key to compositional generalization (Yang et al., 2023). Inspired by stimulus coding within brains (Georgopoulos et al., 1986; O’Keefe & Nadel, 1978; Shadlen & Newsome, 1998), we test various input encodings:
@@
    Ramp coding, an analog ramping code for $(x,y)$ in 2 units.
In each case, the input representation is disentangled w.r.t.  $x$ and $y$ (meaning $x$ and $y$ are independently encoded by separate neurons). These inputs drive an image-generating CNN.
3.2 Model fails to compositionally generalize across various disentangled encodings.
Despite receiving disentangled inputs that correctly specify OOD $(x,y)$ coordinate combinations, the model does not learn to generate the correct outputs beyond the ID region, as seen in the MSE error landscapes of generated images and from the images themselves Fig. 1(b-c). Networks trained with different input encodings, bumps ($x$ and $y$ inputs embedded as the centers of 1D Gaussian bumps in two vectors of length $N$, Fig. 1(b), right) or ramps ($x$ and $y$ represented as analog ramping rates in two units, Fig. 1(c), right) demonstrate different degrees of “local” generalization: some input encodings can support bumps partially inside the OOD region if they overlap with the ID region). However, all networks largely fail in the OOD region, and all consistently generate multiple bumps when prompted with an OOD $(x,y)$ input.
Hence, simply providing factorized inputs, regardless of the form of encodings, does not suffice. The deeper layers appear to re-entangle $x$ and $y$ or rely on memorized local features, leading to poor compositional generalization. While not shown, models trained with disentangled 1-hot and positional input encodings also fail to compositionally generalize. These findings confirm that disentangled latents, across a variety of encoding formats, are not sufficient for compositional generalization.
For the remainder of the manuscript, we use inputs encoded as local bumps or ramps.
@@
The form of failure in OOD generation is suggestive: In many runs, the CNN appears to memorize the training distribution and when prompted with an OOD $(x,y)$ combination, seems to perform a lookup: it finds an image or images in the training data with similar values of $x$ and the closest values of $y$, and superposes these to generate a new image, Fig. 1(b-c), right.
The consistency of these patterns across input encoding formats and architectures indicate a shared underlying mechanism of generalization that we further characterize here.
Figure 2: Cause of failure of compositional generalization despite disentangled inputs/latents: input memorization by decoder undoes factorization. (a-b) show the linear probe metrics of the learned representation as a function of $x$ and $y$ for bump- and ramp-input-encodings, and (c-d) show the factorization score (Eq. 7), and (e-f) show the volume metric $\log_{10}(dv)$, all as a function of layer depth, for models trained with bump-based vs. ramp-based input encodings. The linear probe metric is defined as the $R^{2}$ scores of fitting two linear classifiers with respect to $x$ and $y$.
@@
where $\kappa_{x}\left(x_{i},x_{j}\right)=\kappa_{y}\left(y_{i},y_{j}\right)=\delta_{i,j}$ is given by the Kronecker delta function. With normalization, the Gram matrix element is given by $\tilde{K}_{i,j}=\frac{K_{\text{binary}}(i,j)}{\sum_{j^{\prime}}K_{\text{binary}}\left(i,j^{\prime}\right)}$. Fig. 1(e), we visualize the similarity between representations of OOD-generated samples (sorted by coordinates) and an idealized binary kernel, where similarity equals 1 if coordinates match exactly and 0 otherwise. Specifically, Fig.1(e) compares the ideal binary kernel (top panel) to the actual similarity matrix derived from model-generated Gaussian bump images (bottom panel). The agreement between these two matrices quantifies how closely the model’s learned similarity structure aligns with this ideal binary factorized kernel. Fig. 1(e) corresponds specifically to the top left sections of the full matrices shown in Fig.8 in Appendix F.1, which provide a comprehensive mapping of similarities between all ID and OOD generated samples. These results indicate that the disentangled input structure enables the model to independently map the $x$ and $y$ coordinates to in-distribution (ID) samples, a form of novel OOD generation that is not the same as OOD compositional generalization.
In a set of follow-up experiments on MNIST image rotation, we observe that this same superposition of memorized activation patterns—drawn from related in-distribution data—emerges as a general OOD generalization strategy in neural networks that have learned to memorize ID data. Crucially, this phenomenon applies not only to compositional tasks but also to single‐dimension interpolation and extrapolation. We detail these results in Appendix F.2.
Similarly, in CNN-based guided diffusion models tasked with Gaussian bump image generation, (Liang et al., 2024a) reported that the model generated a superposition of multiple ID bumps when guided to generate an OOD bump. Similar to both these findings, (Kamb & Ganguli, 2024) showed that guided diffusion models generate new images by gluing together local patches from the closest relevant images in the training data. Together with (Lippl & Stachenfeld, 2024), these observations suggest an emerging consensus view of how various generative deep networks function in OOD generation – by combining literal patches of their training data, rather than combining factors from the data, even when provided with information on what these factors should be in the form of disentangled inputs.
@@
Why is disentangled input insufficient for learning disentangled factors of variation? Here we seek deeper mechanistic insight into model’s learned representations along the network layers. Specifically, we track how strongly $x$ and $y$ remain disentangled across layers. While the input layer is disentangled (by construction), subsequent convolutional layers produce mixed features with large cross-terms, particularly near the final decoding layer. This “manifold warping” effectively destroys compositional structure, leaving the network reliant on patterns that do not extend to unseen $(x,y)$ combinations.
@@
$\mathrm{d}v(x,y)=\sqrt{\det\bigl(g(x,y)\bigr{missing})}\mathrm{d}x\,\mathrm{d}y$ (6)
for representation output at each of the network layer. The metric visualizations $\log_{10}(dv)$ are compared between networks with bump-based vs. ramp-based encodings in Fig. 2(e) and (f). Visually, we observe significant distortion as a function of network depth, especially in the OOD region. To accompany the distortion analysis, we plot the linear probe metric (linear classifier $R^{2}$ w.r.t $x$ and $y$) as well as the factorization metric proposed based $g$ in Eq. (7) as a function of network depth in Fig. 2(a-d). Here the factorization metric is defined as follows
@@
where $g_{ij}$’s are the matrix elements of the $2\times 2$ matrix $g$. We note that both metrics decay relatively gracefully over the layer depth for both networks with bump-based and ramp-based input encodings. Intriguingly, shallow networks with ramp-based encoding experiences a sharper drop in factorization. This is likely due to the higher demand for dimension expansion with the ramp-based 2-neuron input, which leads to higher distortion when allowed fewer network layers to accommodate the expansion. The volume metric visualization as well as the linear probe metric as a function of layer depth for different network depths are shown in Fig. 11 and Fig. 12 (Fig. 13 and Fig. 14 for MLPs) as well as generated OOD images in Fig. 10 in Appendix F.3.